5. Potassium chlorate decomposes upon heating as follows:
2KClO3
(s) → 2KCl(s) + 3O2
(g)
A 2.72-g sample of KClO3 decomposes, and the oxygen at 23.4°C and 0.935 atm is
collected. What volume of oxygen gas will be collected, assuming 100% yield?
Answer #1

2 KClO3 (s) ..............>  2 KCl(s) + 3 O2 (g)

2 mole KClO3 decompose to form 2 mole KCl and 3 mole O2.

2.72-g sample of KClO3 = mass / molar mass = 2.72 g / 122.55 g/mole = 0.0222 mole.

thus

mole of O2 produced = (3 / 2) * 0.0222 = 0.0333 mole.

we have

T = (273 + 23.4) = 296.4 K

P = 0.935 atm

n = 0.0333 mole.

Ideal gas equation is

PV = nRT

or

0.935 * V = 0.0333 * 0.0821 * 296.4

or

V = 0.867 L

volume of oxygen gas will be collected = 0.867 L